i need to no how to solve this problem can u show me
2(x+5)+3x=3(-2x-1)

Answer by scott8148(6628)

distributing ___ 2x + 10 + 3x = -6x - 3

collecting terms ___ 5x + 10 = -6x - 3

adding 6x and subtracting 10 ___ 11x = -13

dividing by 11 ___ x = -13/11